# Scaling laws for velocity profile of granular flow in rotating drums

### v2

@@ II.2.1 Time evolution equations
In the DEM simulation [24, 25, 26], we consider $N$ particles in a drum. The dynamics of particle $i$ is given by the time evolution equations for the position ${\bf\it r}_{i}=(X_{i},Z_{i})$ and angular velocity $\omega_{i}$ of particle $i$ as follows:
@@
where $k^{(\mathrm{n})}$ and $\eta^{(\mathrm{n})}_{ij}$ are the normal elastic and the normal viscous constants, respectively; here, $\Theta(x)$ is the Heaviside step function satisfying by $\Theta(x)=1$ for $x\geq 0$, and $\Theta(x)=0$ for otherwise. The normal displacement ${\bf\it\delta}^{(\mathrm{n})}_{ij}$, the normal relative velocity ${\bf\it v}^{(\mathrm{n})}_{ij}$, and the normal unit vector ${\bf\it n}_{ij}$ in Eq. (6) are respectively given by
@@
[…]
@@
The subscript “$\mathrm{stick}$” in Eq. (13) indicates that the integral is performed only when $\left|{\bf\it F}^{(\mathrm{t,st})}_{ij}\right|<\mu_{\mathrm{p}}\left|{\bf\it F}^{(\mathrm{n})}_{ij}\right|$. The torque $T_{ij}$ is given by
@@
II.2.2 Numerical conditions
The diameters of the particles range uniformly from $d$ to $1.2d$ [10]. The mass and moment of inertia of particle $i$ are given by $m_{i}=\frac{1}{4}\pi\rho_{s}d_{i}^{2}$ and $I^{p}_{i}=\frac{1}{8}m_{i}{d_{i}}^{2}$, respectively. We set $k^{(\mathrm{n})}=2.0\times 10^{5}mg/d$, $k^{(\mathrm{t})}/k^{(\mathrm{n})}=2/7$, and $\mu_{p}=0.50$, where $m$ is the mass of the particle with diameter $d$ [27]. The normal viscous constant between particles $i$ and $j$ is given by
$\eta^{(\mathrm{n})}_{ij}=\eta^{(\mathrm{t})}_{ij}=-2\ln{e}\sqrt{\frac{m^{*}_{ij}k^{(\mathrm{n})}}{\pi^{2}+(\ln{e})^{2}}}$ (17)
@@ II.2.2 Numerical conditions
The rough wall surface of the drum is modeled by placing $N_{\rm w}$ particles of diameter $d$ equally spaced on the circle of diameter $D$ to avoid slip at the surface. The particles constituting the rough wall are rotated around the circle with the angular velocity $\Omega$. We set $N_{\rm w}=475$, $632$, and $946$ for the drums with $D=150d,300d$, and $D=450d$, respectively. The drum is half-filled with $N$ particles, where $N$ are set to $7700$, $30600$, and $69200$ for $D=150d,300d$, and $450d$, respectively. Through this numerical simulation, we calculate the velocity distribution in the steady state using the procedure shown in Appendix A.
II.3 CFD simulation
In the CFD simulation, we treat the granular bed in the rotating drum as an incompressible non-Newtonian fluid. The continuum equations for the granular material are numerically solved using the finite difference method with the time step $\Delta t$ and grid width $\Delta x$. The free space in the drum is considered the gas phase to treat the free surface of the granular bed. The time evolution of the two-phase flow with the granular and gas phases is numerically solved by the volume of fluid (VoF) method [28, 29, 30]. This CFD simulation is an extension of the method implemented in Ref. [31] for an incompressible Newtonian fluid in a rotating container.
II.3.1 Governing equations
@@
with
${\bf\it v}_{\mathrm{wall}}=-\Omega Z{\bf\it e}_{X}+\Omega X{\bf\it e}_{Z}.$ (23)
II.3.2 Volume of fluid method
In the VoF method, we introduce the phase volume fraction $\alpha({\bf\it r},t)$ of the fluid phase. Here, $\alpha=1$ and $\alpha=0$ correspond to the areas occupied only by the granular and gas phases, respectively. In the interface between the granular and gas phases, $\alpha$ satisfies $10<\alpha<$ 1. We consider the position with $\alpha=0.5$ as the free surface of the granular bed. The phase volume fraction $\alpha({\bf\it r},t)$ obeys the advection equation
@@ II.3.2 Volume of fluid method
This equation is solved using the WLIC/THINC (weighed line interface calculation/tangent of hyperbola for interface capturing) method [32]. The density $\rho$ and viscosity $\eta$ are calculated as follows:
@@
Here, $\rho_{\mathrm{air}}$ and $\eta_{\mathrm{air}}$ are constants representing the density and viscosity in the gas phase, respectively. The density $\rho_{\mathrm{g}}$ in the granular phase is treated as a constant. The viscosity $\eta_{\mathrm{g}}$ in the granular phase is given by
$\eta_{\mathrm{g}}\left(\dot{\gamma},p\right)=\frac{\mu(I)p}{\dot{\gamma}}$ (27)
in the $\mu(I)$-rheology [19]. We adopt the functional form of $\mu(I)$ as Eq. (2). The inertial number $I$ is given by $I=\dot{\gamma}/\sqrt{p/\rho_{s}d^{2}}$ with
@@
which corresponds to the second invariant of ${\bf\it\dot{\gamma}}$. At constant pressure $p$, the $\mu(I)$-rheology corresponds to the Bingham model.
@@
II.3.3 Simplified marker and cell method
@@
First, for the prediction step, the predicted velocity $\mathbf{v}^{*}$ in ${\bf\it r}\in\mathrm{V_{f}}$ is determined by
@@
where the first term is treated implicitly, and the other terms are treated explicitly [34]. Here, $\nu_{0}$ is the constant component of kinematic viscosity given by
@@
According to Ref. [33], we calculate the pressure $p^{n}$ used in Eq. (29) as the solution of
@@
II.3.4 Boundary data immersion method
We treat the flow driven by the rotating drum using the boundary data immersion (BDI) method [35]. The solid domain $\mathrm{V_{s}}$ is defined as the circular shell with the inner and outer diameters $D$ and $D+16\Delta x$, respectively. In this method, we introduce a single meta-equation to solve the time evolution equations in the fluid and solid domains simultaneously. Here, we define the kernel function
@@ II.3.4 Boundary data immersion method
to continuously connect the fluid and solid domains. Here, the shortest distance from the wall surface at the position ${\bf\it r}$ is denoted by $d_{k}({\bf\it r})$, and $\epsilon$ is a parameter characterizing the thickness of the interface between the fluid and solid domains. From Eq. (41), the kernel function $\chi$ is $1$ in the fluid domain and 00 in the solid domain.
@@
$\nabla\cdot\left\{\frac{\chi}{\rho^{n}}\nabla p^{n}\right\}=\nabla\cdot\left\{\chi\left(-\nabla\cdot\left({\bf\it v}^{n}{\bf\it v}^{n}\right)+\frac{1}{\rho^{n}}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right)+\left(1-\chi\right)\frac{{\bf\it v}_{\mathrm{wall}}}{\Delta t}\right\},$ (46)
@@
${\bf\it v}^{n+1}=\chi\left[{\bf\it v}^{n}+\Delta t\left\{-\nabla\left({\bf\it v}^{n}{\bf\it v}^{n}\right)-\frac{1}{\rho^{n}}\nabla p^{n}+\frac{1}{\rho}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right\}\right]+\left(1-\chi\right){\bf\it v}_{\mathrm{wall}}$ (47)
with Eqs. (20) and (22). We solve Eqs. (45) and (46) using the Bi-CGSTAB method [36].
@@
Note that near the interface between the granular and gas phases, the solution of Eq. (46) may result in pressure below 00, which leads to unphysical behavior when calculating the viscosity of the granular phase $\eta_{\mathrm{g}}$ in Eq. (29) [37, 38]. To prevent this, a sufficiently small positive constant $p_{\mathrm{min}}$ is set as the minimum value of the pressure.
II.3.5 Numerical conditions
@@
We estimate $\mu_{s}$, $\mu_{2}$, and $I_{0}$ in Eq. (2) as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ from the DEM simulation under uniform shear with the parameters in Sec. II.2.2. The details of the estimation are shown in Appendix B. Note that Eq. (2) is also used as the constitutive equation for grains with the Hertzian contact model by choosing appropriate values of the parameters [18]. We set the granular phase density as $\rho_{g}=\rho_{s}\phi$ with $\rho_{s}=m/(\pi d^{2}/4)$. The packing fraction $\phi$ of the particle is set to $0.825$ based on its value in the static flow regime of the rotating drum in the DEM simulation (Appendix C). We use $\eta_{\mathrm{g,max}}=8000m\sqrt{g/d}$, $\rho_{\mathrm{air}}=10^{-3}\rho_{g}$, $\eta_{\mathrm{air}}=10^{-6}\eta_{\mathrm{g,max}}$, and $p_{\mathrm{min}}=10^{-5}mg/d$. We have checked that the velocity profile is quantitatively the same even if we use larger $\eta_{\mathrm{g,max}}$ and lower $\rho_{\mathrm{air}}$, $\eta_{\mathrm{air}}$, and $p_{\mathrm{min}}$. We set $\Delta x=D/75$, $\Delta t=1.0\times 10^{-4}\sqrt{d/g}$, and $\epsilon=2\Delta x$. The maximum error for the Bi-CGSTAB method is set to $e_{\rm max}=10^{-5}$.

### v3

@@ Appendix A Details of DEM simulations
In this Appendix, we present the details of our DEM simulations. The dynamics of particle $i$ is given by the time evolution equations for the position ${\bf\it r}_{i}=(X_{i},Z_{i})$ and angular velocity $\omega_{i}$ of particle $i$ as follows:
@@
where $k^{(\mathrm{n})}$ and $\eta^{(\mathrm{n})}_{ij}$ are the normal elastic and the normal viscous constants, respectively; here, $\Theta(x)$ is the Heaviside step function, defined by $\Theta(x)=1$ for $x\geq 0$ and $\Theta(x)=0$ for $x<0$. The normal displacement ${\bf\it\delta}^{(\mathrm{n})}_{ij}$, the normal relative velocity ${\bf\it v}^{(\mathrm{n})}_{ij}$, and the normal unit vector ${\bf\it n}_{ij}$ in Eq. (24) are respectively given by
@@
$\delta^{(\mathrm{t})}_{ij}=\int_{\mathrm{stick}}dt\ {\bf\it v}^{(\mathrm{t})}_{ij}\cdot{\bf\it t}_{ij},$ (31)
@@
[…]
@@
The diameters of the particles are uniformly distributed from $d$ to $1.2d$. The mass and moment of inertia of particle $i$ are given by $m_{i}=\frac{1}{4}\pi\rho_{s}d_{i}^{2}$ and $I^{p}_{i}=\frac{1}{8}m_{i}{d_{i}}^{2}$, respectively. We use $k^{(\mathrm{n})}=2.0\times 10^{5}mg/d$, $k^{(\mathrm{t})}/k^{(\mathrm{n})}=2/7$, and $\mu_{p}=0.50$. The normal viscous constant between particles $i$ and $j$ is given by
[…]
@@
The rough wall surface of the drum is modeled by placing $N_{\rm w}$ particles of diameter $d$ equally spaced along a circle of diameter $D$ in order to prevent slip at the boundary. These wall particles are rigidly rotated with the angular velocity $\Omega$. We set $N_{\rm w}=475$, $632$, and $946$ for drums with $D=150d$, $300d$, and $450d$, respectively. The drum is half-filled with $N$ particles, where $N=7700$, $30600$, and $69200$ for $D=150d$, $300d$, and $450d$, respectively. To generate the initial packing, the $N$ particles are first placed randomly inside the drum and allowed to settle under gravity without rotation. Subsequently, the drum rotation with angular velocity $\Omega$ is started and the system is evolved until a steady flow state is reached. In the DEM simulations, the packing fraction is not explicitly controlled but is instead determined self-consistently by the particle dynamics under gravity.
Appendix B Velocity field in DEM simulation
The velocity distribution in Fig. 2(a) is calculated using the following procedure in the DEM simulation. First, the system is divided into small square regions with a length of $4.8d$ in the plane of the Cartesian coordinates $\left(X,Z\right)$. Then, the average particle velocity in each region is plotted as the velocity field. The velocity field in the plane of the Cartesian coordinates $(x,z)$ shown in Fig. 3 is numerically obtained as the average particle velocity in the square regions with a length of $1.2d$.
Appendix C Details of CFD simulation
In this Appendix, we present the details of our CFD simulation. In the CFD simulation, we treat the granular bed in the rotating drum as an incompressible non-Newtonian fluid. The continuum equations for the granular material given by Eqs. (1)–(5) are numerically solved using the finite difference method with the time step $\Delta t$ and grid width $\Delta x$. The free space above the granular bed is modeled as a gas phase with very low density and viscosity, allowing the free surface of the granular material to be treated. The time evolution of the two-phase flow with the granular and gas phases is numerically solved by the volume of fluid (VoF) method. [42, 43, 44]
C.1 Governing equations
@@
with ${\bf\it v}_{\mathrm{wall}}$ given by Eq. (7).
C.2 Volume of fluid method
In the VoF method, we introduce the phase volume fraction $\alpha({\bf\it r},t)$ of the fluid phase. Here, $\alpha=1$ and $\alpha=0$ correspond to the areas occupied only by the granular and gas phases, respectively. In the interface between the granular and gas phases, $\alpha$ satisfies $0<\alpha<1$. We consider the position with $\alpha=0.5$ as the free surface of the granular bed. The phase volume fraction $\alpha({\bf\it r},t)$ obeys the advection equation
@@ C.2 Volume of fluid method
This equation is solved using the WLIC/THINC (weighed line interface calculation/tangent of hyperbola for interface capturing) method. [45] The density $\rho$ and viscosity $\eta$ are calculated as follows:
@@
Here, $\rho_{\mathrm{air}}$ and $\eta_{\mathrm{air}}$ are constants representing the density and viscosity in the gas phase, respectively. The density $\rho_{\mathrm{g}}$ in the granular phase is treated as a constant. The viscosity $\eta_{\mathrm{g}}$ in the granular phase is given by Eq. (5) in the $\mu(I)$-rheology. [15] We adopt the functional form of $\mu(I)$ as Eq. (4). The inertial number $I$ is given by $I=\dot{\gamma}/\sqrt{p/(\rho_{s}d^{2})}$ with
@@
which corresponds to the second invariant of $\bf\it\dot{\gamma}$}.
@@
C.3 Simplified marker and cell method
@@
First, for the prediction step, the predicted velocity $\bf\it v$$^{*}$ in ${\bf\it r}\in\mathrm{V_{f}}$ is determined by
@@
where the first term is treated implicitly, and the other terms are treated explicitly. [47] Here, $\nu_{0}$ is the constant component of kinematic viscosity given by
@@
According to Ref. 46, we calculate the pressure $p^{n}$ used in Eq. (45) as the solution of
@@
C.4 Boundary data immersion method
We treat the flow driven by the rotating drum using the boundary data immersion (BDI) method. [48] The solid domain $\mathrm{V_{s}}$ is defined as the circular shell with the inner and outer diameters $D$ and $D+16\Delta x$, respectively. In this method, we introduce a single meta-equation to solve the time evolution equations in the fluid and solid domains simultaneously. Here, we define the kernel function
@@ C.4 Boundary data immersion method
to continuously connect the fluid and solid domains. Here, the shortest distance from the wall surface at the position $\bf\it r$} is denoted by $d_{k}({\bf\it r})$, and $\epsilon$ is a parameter characterizing the thickness of the interface between the fluid and solid domains. From Eq. (57), the kernel function $\chi$ is $1$ in the fluid domain and 0 in the solid domain.
@@
$\nabla\cdot\left\{\frac{\chi}{\rho^{n}}\nabla p^{n}\right\}$ $=$ $\nabla\cdot\left\{\chi\left(-\nabla\cdot\left({\bf\it v}^{n}{\bf\it v}^{n}\right)+\frac{1}{\rho^{n}}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right)\right\}$ (62)
$+\nabla\cdot\left\{\left(1-\chi\right)\frac{{\bf\it v}_{\mathrm{wall}}}{\Delta t}\right\},$
@@
${\bf\it v}^{n+1}$ $=$ $\chi\left[{\bf\it v}^{n}+\Delta t\left\{-\nabla\left({\bf\it v}^{n}{\bf\it v}^{n}\right)-\frac{1}{\rho^{n}}\nabla p^{n}+\frac{1}{\rho}\nabla\cdot{\bf\it\tau}^{n}+{\bf\it g}\right\}\right]$ (63)
$+\left(1-\chi\right){\bf\it v}_{\mathrm{wall}}$
with Eqs. (38) and (40). We solve Eqs. (61) and (62) using the Bi-CGSTAB method. [49]
@@
Note that near the interface between the granular and gas phases, the solution of Eq. (62) may result in pressure below 0, which leads to unphysical behavior when calculating the viscosity of the granular phase $\eta_{\mathrm{g}}$ in Eq. (45). [50, 51] To prevent this, a sufficiently small positive constant $p_{\mathrm{min}}$ is set as the minimum value of the pressure.
C.5 Numerical conditions
@@
We estimate $\mu_{s}$, $\mu_{2}$, and $I_{0}$ in Eq. (4) as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ from the DEM simulation under uniform shear with the parameters in Sec. II.1. The details of the estimation are shown in Appendix D. We set the granular density as $\rho_{g}=\rho_{s}\phi$ with $\rho_{s}=m/(\pi d^{2}/4)$. The packing fraction $\phi$ of the particle is set to $0.825$ based on its value in the static flow regime of the rotating drum in the DEM simulation (Appendix E). We use $\eta_{\mathrm{g,max}}=8000m\sqrt{g/d}$, $\rho_{\mathrm{air}}=10^{-3}\rho_{g}$, $\eta_{\mathrm{air}}=10^{-6}\eta_{\mathrm{g,max}}$, and $p_{\mathrm{min}}=10^{-5}mg/d$. We have checked that the velocity profile is quantitatively the same even if we use larger $\eta_{\mathrm{g,max}}$ and lower $\rho_{\mathrm{air}}$, $\eta_{\mathrm{air}}$, and $p_{\mathrm{min}}$. We set $\Delta x=D/75$, $\Delta t=1.0\times 10^{-4}\sqrt{d/g}$, and $\epsilon=2\Delta x$. The maximum error for the Bi-CGSTAB method is set to $e_{\rm max}=10^{-5}$.